A boy puts a heavy box of mass M on his head and jumps down from the top of a multi-storeyed
building to the ground. How much is the force exerted by the box on his head during his free fall?
Does the force greatly increase during the period he balances himself after striking the ground?

Grade:11

1 Answers

Arun
6 years ago
both boy and box M are falling freely from building , at building force exerted by box on boy is due to its wt. on falling aap wt of box is = m( g - a)= m( g - g ) ( g = a )= 0so net force is 0 , the force doesnot increase